# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 6 Non-simply connected cases
$SN^{*}(\partial V)$ is then identified with the normal bundle $N_{\partial V/\partial W_{0}}$ of $\partial V\subset\partial W_{0}$ upon the attaching of $\partial_{-}\textsf{H}_{V}=\partial V\times D^{2}_{y}$ to $\partial W_{0}$ along $\partial V$. As we go once along $\partial V$ the contact framing $\nabla h$ makes one positive full rotation relative to the canonical framing $\partial_{y_{1}}$. In other words, the canonical framing $\partial_{y_{1}}$ along $\partial V$ is obtained from the contact framing by adding one left twist. So the natural framing along $\partial V$ given by $\partial_{y_{1}}$ is $-1$ relative to the contact framing along $\partial V$, which meets with the Stein condition on the attachment of 4-dimensional 2-handles.
The above result on the framing criteria along $\partial V\subset\partial_{-}\textsf{H}$ apply to $V_{i}$ for all $i=1,2,\dots,n$ as well. Each of the 2-handlebodies $\textsf{H}_{V_{i}}$ has its core disc $V_{i}$ attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist. In particular, $W_{n}=\EuFrak{r}^{-1}(D)=\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})$ admits a Stein structure by Theorem 6.1.
Case 2: $\textsf{H}_{C_{j}}$.  Pick any one of the 2-dimensional handles $H_{C_{j}}$ and denote it as $H_{C}$, where $C$ is the core curve of $H_{C}$. We will see that $\EuFrak{r}^{-1}(H_{C})$ can be identified with the union of a 4-dimensional 1-handle $\textsf{H}_{\gamma}$ with core curve $\gamma\subset h^{-1}(0)$ and a 4-dimensional 2-handle $\textsf{H}_{U}$ with core disc $U\subset\cap h^{-1}(0)$, and $\partial U$ is attached to $\partial\textsf{H}_{\gamma}\cap h^{-1}(0)$.
The preimage $L:=\EuFrak{r}^{-1}_{0}(C)\subset h^{-1}(0)\cap W_{n}$ is a Lagrangian annulus $L\cong I\times S^{1}$ foliated by $G$-orbits generated by the Hamiltonian vector field $X$. We may parametrize $L$ as
@@
so that $X=\partial_{x_{2}}$, and each integral curve of $\partial_{x_{1}}$ is a lifting of $C$ in $L$. The boundary of $L$, which is the pair of $G$-orbits given by the equations $x_{1}=\pm 1$ is attached to $\partial W_{n}$. Let
@@
for some $0<\epsilon\ll 1$, with $\partial_{-}\textsf{H}_{\gamma}:=\partial\gamma\times I^{3}$ attached to $\partial M_{0}$. Here $y_{2}$ parametrizes integral curves of $-\nabla h$, the negative gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metric on $M$. Let
@@
$\gamma_{-}$ =$=\{-1\leq x_{1}\leq 1,\ x_{2}=\epsilon\},$ $\quad\dot{\gamma}_{-}$ $=\partial_{x_{1}},$ (9)
$\sigma_{+}$ =$=\{x_{1}=1,\ \epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{+}$ $=\partial_{x_{2}}=X,$ (10)
$\gamma_{+}$ =$=\{-1\leq x_{1}\leq 1,\ x_{2}=2\pi-\epsilon\},$ $\quad\dot{\gamma}_{+}$ $=-\partial_{x_{1}},$ (11)
$\sigma_{-}$ =$=\{x_{1}=-1,\ -\epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{-}$ $=-\partial_{x_{2}}=-X.$ (12)
Recall the $G$-invariant primitive 1-form $\lambda$ of $\omega$. Observe that $\EuFrak{r}^{-1}(\partial D)\subset\partial W_{n}$ is a subdomain of $\partial W_{n}$, $\lambda$ restricted to $\partial W_{n}$ is a contact 1-form near $\partial L$, whose contact structure along $\partial L$ is spanned by the symplectic pair $\nabla h$ and $X$. So along $\sigma_{+}$ the contact structure is $X\wedge(-\nabla h)=\partial_{x_{2}}\wedge\partial_{y_{2}}$, and along $\sigma_{-}$ the contact structure is $-X\wedge\nabla h=-\partial_{x_{2}}\wedge(-\partial_{y_{2}})$. Along the boundary $(\partial_{+}\textsf{H}_{\gamma})\cap L=\gamma_{-}\cup\gamma_{+}$ the vector field $X$ is normal to $\partial_{+}\textsf{H}_{\gamma}$, hence along $\gamma_{-}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $\partial_{x_{1}}\wedge\partial_{y_{1}}$, spanned by the symplectic pair $\dot{\gamma}_{-}=\partial_{x_{1}}$ and $\partial_{y_{1}}$. Similarly along $\gamma_{+}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $(-\partial_{x_{1}})\wedge(-\partial_{y_{1}})$, spanned by the symplectic pair $\dot{\gamma}_{+}=-\partial_{x_{1}}$ and $-\partial_{y_{1}}$.
@@
˙$\begin{split}\dot{\ell}(\theta)&=\cos\theta\partial_{x_{2}}-\sin\theta\partial_{x_{1}},\quad\ell(0)=(1,0)\\ &=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}},\end{split}$
@@
Similar to Case 1, the framing of the symplectic normal bundle $SN^{*}(\partial U)$ (which is spanned by the basis vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$) is given by $\partial_{y_{1}}$. As we go once along $\ell$ the contact framing makes one positive full rotation relative to $\partial_{y_{1}}$. So the canonical framing $\partial_{y_{1}}$ is $-1$ relative to the contact framing along $\ell$. This result applies to all $\textsf{H}_{C_{i}}$. Combining the result from Case 1 we conclude that $M$ admits a Stein structure following Theorem 6.1. This completes the proof of Theorem 1.7.

### v4

@@ 6 Non-simply connected cases
$SN^{*}(\partial V)$ is then identified with the normal bundle $N_{\partial V/\partial W_{0}}$ of $\partial V\subset\partial W_{0}$ upon the attaching of $\partial_{-}\textsf{H}_{V}=\partial V\times D^{2}_{y}$ to $\partial W_{0}$ along $\partial V$. As we go once along $\partial V$ the contact framing $\nabla h$ makes one positive full rotation relative to the canonical framing $\partial_{y_{1}}$. In other words, the canonical framing $\partial_{y_{1}}$ along $\partial V$ is obtained from the contact framing by adding one left twist. So the natural framing along $\partial V$ given by $\partial_{y_{1}}$ is 1-1 relative to the contact framing along $\partial V$, which meets with the Stein condition on the attachment of 4-dimensional 2-handles.
The above result on the framing criteria along $\partial V\subset\partial_{-}\textsf{H}$ apply to $V_{i}$ for all $i=1,2,\dots,n$ as well. Each of the 2-handlebodies $\textsf{H}_{V_{i}}$ has its core disc $V_{i}$ attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist. In particular, $\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})$ admits a Stein structure by Theorem 6.1. $\Box$
Case 2: $\textsf{H}_{C_{j}}$. Pick any one of the 2-dimensional handles $H_{C_{j}}$ and denote it as $H_{C}$, where $C$ is the core curve of $H_{C}$. We will see that $\EuFrak{r}^{-1}(H_{C})$ can be identified with the union of a 4-dimensional 1-handle $\textsf{H}_{\gamma}$ with core curve $\gamma\subset h^{-1}(0)$ and a 4-dimensional 2-handle $\textsf{H}_{U}$ with core disc $U\subset h^{-1}(0)$, and $\partial U$ is attached to $\partial\textsf{H}_{\gamma}\cap h^{-1}(0)$.
The preimage $L:=\EuFrak{r}^{-1}_{0}(C)\subset h^{-1}(0)$ is a Lagrangian annulus $L\cong I\times S^{1}$ foliated by $G$-orbits generated by the Hamiltonian vector field $X$. We may parametrize $L$ as
@@
so that $X=\partial_{x_{2}}$, and each integral curve of $\partial_{x_{1}}$ is a lifting of $C$ in $L$. The boundary of $L$, which is the pair of $\Phi$-orbits given by the equations $x_{1}=\pm 1$ is attached to $\partial W_{n}$. Let
@@
for some $0<\epsilon\ll 1$, with $\partial_{-}\textsf{H}_{\gamma}:=\partial\gamma\times I^{3}$ attached to $\partial M_{0}$. Here $y_{2}$ parametrizes integral curves of $\nabla h$, the negative gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metric on $M$. Let
@@
(8) $\gamma_{-}$ $=\{-1\leq x_{1}\leq 1,\ x_{2}=\epsilon\},$ $\quad\dot{\gamma}_{-}$ $=\partial_{x_{1}},$
(9) $\sigma_{+}$ $=\{x_{1}=1,\ \epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{+}$ $=\partial_{x_{2}}=X,$
(10) $\gamma_{+}$ $=\{-1\leq x_{1}\leq 1,\ x_{2}=2\pi-\epsilon\},$ $\quad\dot{\gamma}_{+}$ $=-\partial_{x_{1}},$
(11) $\sigma_{-}$ $=\{x_{1}=-1,\ -\epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{-}$ $=-\partial_{x_{2}}=-X.$
Recall the $\Phi$-invariant primitive 1-form $\lambda$ of $\omega$. Observe that $\EuFrak{r}^{-1}(\partial D)\subset\partial W_{n}$ is a subdomain of $\partial W_{n}$, $\lambda$ restricted to $\partial W_{n}$ is a contact 1-form near $\partial L$, whose contact structure along $\partial L$ is spanned by the symplectic pair $\nabla h$ and $X$. So along $\sigma_{+}$ the contact structure is $X\wedge(-\nabla h)=\partial_{x_{2}}\wedge\partial_{y_{2}}$, and along $\sigma_{-}$ the contact structure is $X\wedge\nabla h=-\partial_{x_{2}}\wedge(-\partial_{y_{2}})$. Along the boundary $(\partial_{+}\textsf{H}_{\gamma})\cap L=\gamma_{-}\cup\gamma_{+}$ the vector field $X$ is normal to $\partial_{+}\textsf{H}_{\gamma}$, hence along $\gamma_{-}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $\partial_{x_{1}}\wedge\partial_{y_{1}}$, spanned by the symplectic pair $\dot{\gamma}_{-}=\partial_{x_{1}}$ and $\partial_{y_{1}}$. Similarly along $\gamma_{+}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $(-\partial_{x_{1}})\wedge(-\partial_{y_{1}})$, spanned by the symplectic pair $\dot{\gamma}_{+}=-\partial_{x_{1}}$ and $\partial_{y_{1}}$.
@@
$\begin{split}\dot{\ell}(\theta)&=\cos\theta\partial_{x_{2}}-\sin\theta\partial_{x_{1}},\quad\ell(0)=(1,0)\\ &=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}},\end{split}$
@@
Similar to Case 1, the framing of the symplectic normal bundle $SN^{*}(\partial U)$ (which is spanned by the basis vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$) is given by $\partial_{y_{1}}$. As we go once along $\ell$ the contact framing makes one positive full rotation relative to $\partial_{y_{1}}$. So the canonical framing $\partial_{y_{1}}$ is 1-1 relative to the contact framing along $\ell$. This result applies to all $\textsf{H}_{C_{i}}$. Combining the result from Case 1 we conclude that $M$ admits a Stein structure following Theorem 6.1. This completes the proof of Theorem 1.11. $\Box$